OpenStudy (anonymous):

Find the x-intercepts of the graph of the equation: y = x2 - 3x + 2

OpenStudy (anonymous):

First, you have to factor the equation

OpenStudy (anonymous):

How would i do that?

OpenStudy (anonymous):

i plugged in 0 for y

OpenStudy (anonymous):

If you can't factor polynomials, you need to learn that before proceeding with this. It's a pretty broad topic that's outside of the scope of this.

OpenStudy (anonymous):

I dont even know what a polynomials is... im only in algabra 1...

OpenStudy (anonymous):

That really should've been taught before this problem was given to you.

OpenStudy (anonymous):

One of the many cons of homeschooling... you teach yourself..

OpenStudy (anonymous):

Ah. I suppose I can use this as an example to help you with factoring. You know the FOIL method, right?

OpenStudy (anonymous):

Sadly No.

OpenStudy (anonymous):

I think you're a bit ahead of yourself in your schooling then. FOIL should've been taught first, as a way to simplify things like (x - 1)(x - 1), and then factoring should've been taught as a way to change things like what you've got back in to that form

OpenStudy (anonymous):

I i know the foil Method I just didn't give it a name neither did the book that taught me it...

OpenStudy (anonymous):

Ah, then factoring is a way to change \[y = x^2 + 3x + 2\] back in to that other form. Basically you're looking for the two binomials that, when multiplied, equal your expression

OpenStudy (anonymous):

I know that, But I tried 1-10 Replacing X And they All Didnt Work Could It Be A Decimal or Fraction?

OpenStudy (anonymous):

It isn't, but the numbers in the two binomials can be different, and they can be positive or negative. A few things to help you is that the product of the two numbers in the polynomial have to equal the last term in your equation. I'll give you a hint in that both of your numbers in your binomials are negative, and they're different.

OpenStudy (anonymous):

Are They Withing The -1 To -10 Range?

OpenStudy (anonymous):

I'll do you one better, they're within the -1 to -5 range. Basically, they should look like (x - something)(x - something else) okay? :D

OpenStudy (anonymous):

-3 And -5?

OpenStudy (anonymous):

Not quite. (x - 3)(x - 5) equals out to\[x^2 - 8x + 15\] since this is your first one I'll tell you one of the terms, it'll be (x - 1)(x - something).

OpenStudy (anonymous):

(x - 1)(x - 2?)

OpenStudy (anonymous):

or (x - 1)(x - 4?)

OpenStudy (anonymous):

Yes to the first :D. You can get the x intercepts by setting each of them to 0 individually, like x - 1 = 0 and the same for the other.

OpenStudy (anonymous):

oh ok i think i get it?

OpenStudy (anonymous):

Thanks for your Help!

OpenStudy (anonymous):

Good luck on your future studies!

OpenStudy (anonymous):

So is the Answer x = -1, -2?

OpenStudy (anonymous):

No, it's x = 1, x = 2. 0 = x - 1 Add one to each side 1 = x 0 = x - 2 Add two to each side 2 = x

OpenStudy (anonymous):

Ok Just Clarifying Thanks!